# Differentially Private Matchings

### v1

@@ 5.3 Utility and Number of Rounds
We now show that if all edges become unhopeful, then all nodes are either matched or all of their neighbors are matched. An edge becomes unhopeful if at least one of its endpoints satisfies the matching condition. By Lemma 2.18, a node satisfies the matching condition if and only if $|M(u)|\geq 1$ with probability at least $1-\frac{1}{n^{2c}}$. Hence, if a node is adjacent to all unhopeful edges, then it either satisfies the matching condition and is matched to at least one neighbor with high probability, or it is adjacent to endpoints which are matched with high probability.
@@
Finally, by Lemma 2.18, $|M(u)|\geq 2\cdot 147c\log(n)/\varepsilon^{\prime}$ with probability at most $\frac{1}{n^{2c}}$. By union bound over all probabilities of failure, with probability at least $1-\frac{1}{n^{c}}$ for any constant $c\geq 1$, we obtain a maximal matching, with high probability, in $O(\log n)$ rounds when $b=\Omega(\log^{2}(n)/\varepsilon)$.
@@
Combining Lemma 5.1 and Lemma 5.5 yields the proof of Algorithm 3.
6 Node Differentially Private Matchings
Using sparsification techniques, we also demonstrate the first connection between sparsification and node-differentially private algorithms via arboricity. The class of bounded arboricity [^8] Arboricity is defined as the minimum number of forests to decompose the edges in a graph. A $n$ degree star has max degree $n-1$ and arboricity $1$. graphs is a more general class of graphs than bounded degree graphs; a simple example of a graph with large degree but small arboricity is a collection of stars. For simplicity of presentation, we first present a result assuming a public bound $\tilde{\alpha}$. Then our algorithm is always private but the approximation guarantees hold when $\tilde{\alpha}$ upper bounds the arboricity of the input graph. When such a public bound is unavailable, we show that guessing the bound with powers of 2 achieves the same guarantees up to logarithmic factors.
@@ 6.1 Bounded Arboricity Sparsifiers
Let $\pi\in P_{\binom{n}{2}}$ be a total ordering over unordered vertex pairs and $G\sim G^{\prime}$ be node neighboring graphs. Then the edge edit distance between $H:=\textsc{ContractionSparsify}_{\pi}(G,\Lambda)$ and $H^{\prime}:=\textsc{ContractionSparsify}_{\pi}(G^{\prime},\Lambda)$ is at most $2\Lambda$.
@@
The original sparsification algorithm in [solomon2021local] marks an arbitrary set of $\Lambda$ edges incident to every vertex and takes the subgraph consisting of all edges marked by both endpoints. In our setting, $\pi$ determines the arbitrary marking in our graphs. Hence, our sparsification procedure satisfies the below guarantee.
@@
Let $\pi\in P_{\binom{n}{2}}$ be a total order over unordered vertex pairs, $G$ be a graph of arboricity at most $\alpha$, $\Lambda:=5(1+\nicefrac{{5}}{{\eta}})\cdot 2\alpha$ for some $\eta\in(0,1]$, and $H=\textsc{ContractionSparsify}_{\pi}(G,\Lambda)$. Then if $\mu(\cdot)$ denotes the size of a maximum matching of the input graph,
@@
In particular, any $(\beta,\zeta)$-approximate maximum matching for $H$ is an $(\beta(1+\eta),\zeta(1+\eta))$-approximate matching of $G$.
6.2 Node-DP Maximum Matching
In this section, we design a node-DP algorithm to output an implicit matching in the central model using the sparsification techniques derived in Section 6.1 and any edge DP algorithm (e.g. Corollary 1.3) that outputs implicit solutions in the $\varepsilon$-DP setting.
The algorithm first discards edges according to Theorem 6.2 until there are at most $\Lambda=O(\tilde{\alpha}/\eta)$ edges incident to each vertex. This ensure that the edge-edit distance of node-neighboring graphs is at most $2\Lambda$ (Proposition 6.1). We can thus run any $(\frac{\varepsilon}{2\Lambda})$-edge DP algorithm to ensure $\varepsilon$-node DP.
@@ 6.2 Node-DP Maximum Matching
Fix $\eta\in(0,1]$. Given a public bound $\tilde{\alpha}$ on the arboricity $\alpha$ of the input graph, there is an $\varepsilon$-node DP algorithm that outputs an implicit $b$-matching. Moreover, with probability at least $1-1/\operatorname{poly}(n)$, (i) $b=O\left(\frac{\tilde{\alpha}\log(n)}{\eta\varepsilon}\right)$and (ii) if $\tilde{\alpha}\geq\alpha$, the implicit solution contains a $(2+\eta)$-approximate maximum matching.
@@
6.3 Removing the Assumption on Public Bound
In practice, it is not always possible to obtain a good public bound on the arboricity. Thus we show that this assumption can be removed at the cost of $O\left(\frac{\log^{2}(n)}{\varepsilon}\right)$ additive error. At a high level, we run the conditional node-DP algorithm (Theorem 6.3) with $\tilde{\alpha}=2^{k}$ for $k=1,\dots,\lceil\log_{2}(n)\rceil$ and output the best solution. A key subroutine we need to implement is how to privately check the quality of an implicit solution.
@@ 6.3 Removing the Assumption on Public Bound
Given a fixed implicit $b$-matching, there is an $\varepsilon$-node DP algorithm that estimates the size of the largest matching in the subgraph induced by the implicit $b$-matching with pure additive error $O(\log(n)/\varepsilon)$, with probability $1-1/\operatorname{poly}(n)$.
@@
For a fixed implicit $b$-matching, two node-neighboring graphs induce two node-neighboring subgraphs. The sensitivity of the maximum matching cardinality is $1$ on node-neighboring graphs. Thus computing the exact maximum matching size in the induced subgraph and adding Laplace noise is $\varepsilon$-node DP. ∎
We are now ready to remove the assumption on the public bound $\tilde{\alpha}$ and prove Theorem 1.7, which we restate below for convenience.
See 1.7
Proof.
Our algorithm first computes an $(\varepsilon/3)$-node DP $(1,O(\log(n)/\varepsilon))$-approximate estimate $\hat{\mu}$ on the size of the maximum matching $\mu$ in the input graph (Lemma 6.4). Then, we run the conditional node-DP algorithm (Theorem 6.3) with public bound $\tilde{\alpha}=2^{k}$ for $k=1,\dots,\lceil\log_{2}(n)\rceil$ to obtain $(\varepsilon/3\log(n))$-node DP implicit $b_{k}$-matchings for $k=1,\dots,\lceil\log_{2}(n)\rceil$. We also compute $(\varepsilon/3\log(n))$-node DP estimates $\hat{\mu}_{k}$ of the size of the maximum matching contained within the implicit $b_{k}$-matching (Lemma 6.4). Remark that $\hat{\mu}_{k}$ is a $\left(1,O\left(\log^{2}(n)/\varepsilon\right)\right)$-approximate estimate. Finally, choose $\bar{k}$ to be the smallest $k$ such that
$\hat{\mu}_{k}\geq\frac{1}{2+\eta}\hat{\mu}-O\left(\frac{\log^{2}(n)}{\varepsilon}\right)\,.$
The final output is the $\bar{k}$-th implicit $b_{\bar{k}}$-matching.
The privacy guarantee follows by simple composition. Let $k^{\star}$ be such that $2^{k^{\star}-1}<\alpha\leq 2^{k^{\star}}$. The utility guarantee follows since the $k^{\star}$-th $b_{k^{\star}}$-matching satisfies $b_{k^{\star}}=O\left(\frac{\alpha\log^{2}(n)}{\eta\varepsilon}\right)$ and contains a $(2+\eta)$-approximate maximum matching with probability $1-1/\operatorname{poly}(n)$. ∎
We note that we can remove a factor of $O(\log(n))$ for both the value of $b$ and the additive error from Theorem 1.7 using private selection [cohen2022generalized]. This overhead is a result of privately “selecting” the best hyperparameter and can be reduced to a constant at the cost of additional polynomial computation time.
7 Matchings in the Continual Release Model
@@
We now give algorithms for maximal matching in the continual release model. We give algorithms that satisfy edge-DP and node-DP in two different types of input streams. We consider two types of insertion-only streams where nodes and edges can be inserted but not deleted: arbitrary-order edge arrival streams and adjacency-list order streams.
In the arbitrary-order edge arrival stream, edge insertions arrive in an arbitrary order. In the adjacency-list order insertion streams, nodes arrive in an arbitrary order and once a node arrives, all edges adjacent to the node arrive in an arbitrary order. In the adjacency-list order stream, all edges arrive twice, once per endpoint in the stream.
Before proceeding with our results, we state a variant of the 1-dimensional SVT which allows us to answer “above” $c$ times.
7.1 Multi-Response Sparse Vector Technique
[…]
Theorem 7.1 (Theorem 2 in [lyu2017understanding]).
Algorithm 5 is $\varepsilon$-DP.
2 Class SVT($\varepsilon,\Delta,c$)
3       $\varepsilon_{1},\varepsilon_{2}\leftarrow\nicefrac{{\varepsilon}}{{2}}$
4       $\rho\leftarrow\text{Lap}(\nicefrac{{\Delta}}{{\varepsilon_{1}}})$
5       $\operatorname{count}\leftarrow 0$
6
7       Function ProcessQuery($f_{t}(D),\tau_{t}$)
8             if $\operatorname{count}>c$ then
9                   return “abort”
10
12                   return “above”
13                   $\operatorname{count}\leftarrow\operatorname{count}+1$
14
15            else
16                   return “below”
17
18
19
Algorithm 5 Sparse Vector Technique
7.2 Arbitrary Edge-Order Streams
We now show that our edge and node-DP implicit matching algorithms can be implemented in the arbitrary edge-order continual release model. To avoid redundancy, we present the edge-DP algorithm and then briefly discuss the minor changes to obtain a node-DP algorithm.
See 1.9
Theorem 7.2.
Fix $\eta\in(0,1]$. Given a public bound $\tilde{\alpha}$ on the arboricity $\alpha$ of the input graph, there is an $\varepsilon$-node DP algorithm in the arbitrary order continual release model that outputs implicit $b$-matchings. Moreover, with probability at least $1-1/\operatorname{poly}(n)$, (i) $b=O\left(\frac{\tilde{\alpha}\log^{2}(n)}{\eta^{2}\varepsilon}\right)$and (ii) each implicit solution contains a $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate maximum matching.
We can similarly remove the dependence on a public bound of the node-DP algorithm by guessing the arboricity of the graph to yield Theorem 1.10, our main continual release node-DP result restated below. See 1.10 Similar to Theorem 1.7, we can shave off a factor of $O(\log(n))$ in both the value of $b$ and the additive error from Theorem 1.10 using private selection [cohen2022generalized].
@@ 7.2 Arbitrary Edge-Order Streams
We begin with the pseudocode of $\varepsilon$-edge DP algorithm in Algorithm 6. Then, we prove privacy and utility separately. The proof of privacy is non-trivial since we cannot simply apply composition. Instead, we directly argue by the definition of privacy.
@@ 7             $j_{t}\leftarrow j_{t}+1$
10             $\operatorname{solution}\leftarrow$ ($\nicefrac{{\varepsilon}}{{3c}}$)-edge DP implicit $b$-matching using Algorithm 2 on input $G_{t}$ (Corollary 1.3)
11             $\operatorname{estimate}\leftarrow$ ($\nicefrac{{\varepsilon}}{{3c}}$)-edge DP estimate of the size of largest matching contained within solution (Lemma 6.4)
@@ 13       Output solution, estimate
Now, conditioned on the event that $j_{t-1}=J_{t-1}$, it holds that $j_{t}$ is independent of $\operatorname{solution}_{t-1}$ and $\operatorname{estimate}_{t-1}$. Hence the first product is at most $e^{\varepsilon/3}$ by the privacy guarantees of the SVT (Theorem 7.1). For the second and third products, remark that given $j_{t}=J_{t-1}=\tilde{j}_{t}$, then $\operatorname{solution}_{t}=B_{t-1}=\widetilde{\operatorname{solution}}_{t}$ with probability $1$ and similarly for $\operatorname{estimate}_{t}$. Since $(J_{t})_{t}$ is a non-decreasing non-negative integer sequence bounded above by $c=a\log(n)/\eta$, at most $c=a\log(n)/\eta$ of the ratios from the second and third products are not 1. We can bound each of these ratios by $e^{\varepsilon/3c}$ using the individual privacy guarantees of Algorithm 2 (Corollary 1.3) and the Laplace mechanism. This concludes the proof. ∎

### v2

@@ 5.2.2 Utility and Number of Rounds
We now show that if all edges become unhopeful, then all nodes are either matched or all of their neighbors are matched. An edge becomes unhopeful if at least one of its endpoints satisfies the matching condition. By Lemma A.1, a node satisfies the matching condition if and only if $|M(u)|\geq 1$ with probability at least $1-\frac{1}{n^{2c}}$. Hence, if a node is adjacent to all unhopeful edges, then it either satisfies the matching condition and is matched to at least one neighbor with high probability, or it is adjacent to endpoints which are matched with high probability.
@@
Finally, by Lemma A.1, $|M(u)|\geq 2\cdot 147c\log(n)/\varepsilon^{\prime}$ with probability at most $\frac{1}{n^{2c}}$. By Lemma 4.2 and a union bound over all probabilities of failure, with probability at least $1-\frac{1}{n^{c}}$ for any constant $c\geq 1$, we output a $b$-matching that is at least $\frac{1}{2}$ the size of a maximum matching in $O(\log n)$ rounds when $b=\Omega(\log^{2}(n)/\varepsilon)$.
@@
Combining Lemma 5.1 and Lemma 5.2 yields the proof of Theorem 1.6.
6 Node Differentially Private Implicit Matchings
Using sparsification techniques, we demonstrate the first connection between sparsification and node-differentially private algorithms via arboricity. In particular, we prove Theorem 1.8, which we restate below. See 1.8
The class of bounded arboricity [^8] Arboricity is defined as the minimum number of forests to decompose the edges in a graph. A $n$ degree star has max degree $n-1$ and arboricity $1$. graphs is a more general class of graphs than bounded degree graphs; a simple example of a graph with large degree but small arboricity is a collection of stars. For simplicity of presentation, we first design an algorithm assuming a public bound $\tilde{\alpha}$. Then our algorithm is always private but the approximation guarantees hold when $\tilde{\alpha}$ upper bounds the arboricity of the input graph. When such a public bound is unavailable, we show that guessing the bound with powers of 2 achieves the same guarantees up to logarithmic factors.
@@ 6.1 Bounded Arboricity Sparsifiers
Let $\pi\in P_{\binom{n}{2}}$ be a total ordering over unordered vertex pairs and $G\sim G^{\prime}$ be node neighboring graphs. Then the edge edit distance between $H\coloneqq\textsc{ContractionSparsify}_{\pi}(G,\Lambda)$ and $H^{\prime}\coloneqq\textsc{ContractionSparsify}_{\pi}(G^{\prime},\Lambda)$ is at most $2\Lambda$.
@@
The original sparsification algorithm in [solomon2021local] marks an arbitrary set of $\min(\deg(v),\Lambda)$ edges incident to each vertex $v$ and takes the subgraph consisting of all edges marked by both endpoints. In our setting, $\pi$ determines the arbitrary marking in our graphs. Hence, our sparsification procedure satisfies the below guarantee.
@@
Let $G$ be a graph of arboricity at most $\alpha$ and $\Lambda\coloneqq 5(1+\nicefrac{{5}}{{\eta}})\cdot 2\alpha$ for some $\eta\in(0,1]$. Suppose $H$ is obtained by marking an arbitrary set of $\min(\deg(v),\Lambda)$ edges incident to each vertex $v$ and taking the subgraph consisting of all edges marked by both endpoints. Then if $\mu(\cdot)$ denotes the size of a maximum matching of the input graph,
@@
In particular, any $(\beta,\zeta)$-approximate maximum matching for $H$ is an $(\beta(1+\eta),\zeta(1+\eta))$-approximate matching of $G$. Moreover, this holds for $H=\textsc{ContractionSparsify}_{\pi}(G,\Lambda)$ where $\pi\in P_{\binom{n}{2}}$ is a total order over unordered vertex pairs.
6.2 Proof of Theorem 1.8
6.2.1 Node-DP Maximum Matching with Public Arboricity Bound
In this section, we design a node-DP algorithm to output an implicit matching in the central model using the sparsification techniques derived in Section 6.1 and any edge DP algorithm (e.g. Theorem 1.3) that outputs implicit solutions in the $\varepsilon$-DP setting.
@@ 6.2.1 Node-DP Maximum Matching with Public Arboricity Bound
Fix $\eta\in(0,1]$. Given a public bound $\tilde{\alpha}$ on the arboricity $\alpha$ of the input graph, there is an $\varepsilon$-node DP algorithm that outputs an implicit $b$-matching. Moreover, with probability at least $1-1/\operatorname{poly}(n)$, (i) $b=O\left(\frac{\tilde{\alpha}\log(n)}{\eta\varepsilon}\right)$and (ii) if $\tilde{\alpha}\geq\alpha$, the implicit solution is at least $\frac{1}{2+\eta}$ the size of a maximum matching.
@@
Proof Sketch.
The algorithm first discards edges according to Theorem 6.2 until there are at most $\Lambda=O(\tilde{\alpha}/\eta)$ edges incident to each vertex. This ensure that the edge-edit distance of node-neighboring graphs is at most $2\Lambda$ (Proposition 6.1). We can thus run any $(\frac{\varepsilon}{2\Lambda})$-edge DP algorithm to ensure $\varepsilon$-node DP. ∎
6.2.2 Removing the Assumption of Public Bound
In practice, it is not always possible to obtain a good public bound on the arboricity. However, we can always compute an $\varepsilon$-node DP estimate of the arboricity up to $O(\log(n)/\varepsilon)$ additive error since the sensitivity of the arboricity is $1$ even for node-neighboring graphs. This incurs a slight blow-up in $b$ and accounting for this blowup leads to a proof of Theorem 1.8.
7 Improved Node-DP Implicit Bipartite Matchings
In this section, we restate and prove Theorem 1.10. See 1.10
The main difference in our algorithm is in implementing the counters. In the private algorithm of [hsu2014private], each of these counters need to be accurate at each of the $nT$ iterations. This causes significant privacy loss. Our algorithm has a different implementation (of the same basic algorithm), where we use the fact that only the counts at the end of each of the $T$ iterations are important. Moreover, our implementation does not need the counts themselves, but just needs to check whether they are above the supply $s$. This allows us to use the Multidimensional AboveThreshold Mechanism (Algorithm 1) to get further improvements.
3 for $t\leftarrow 1$ to $T$ do
4       foreach item $u\in V_{L}$ do
6
      // Each bidder $v$ proposes to the neighbor $u$ with lowest price
7       foreach bidder $v\in V_{R}$ do
8             if $v$ is not matched then
10                   if $p_{u}\leq 1$ then
11                         Denote bidder $v$ as (temporarily) matched with item $u$ for timestep $t$
12                         $c_{u}\leftarrow c_{u}+1$
13
14
15
16      foreach item $u\in V_{L}$ do
17             Release $p_{u}$
19                   $p_{u}\leftarrow p_{u}+\eta$
21
22
23       foreach item $u\in V_{L}$ do
24             Release startu
27             while $M_{u}+\mathrm{Lap}(4/\varepsilon^{\prime})\geq\tilde{s}_{u}$ do
28                   if a bidder $v$ matched with item $u$ at timestep $\text{start}_{u}$ then
29                         Mark bidder $v$ as unmatched
30
32
33
      // Terminate early if the algorithm is close to convergence
34       Let $c_{0}$ denote the number of bids made in this iteration
36             Terminate the algorithm
37
38
7.1 Proof of Theorem 1.10
The pseudocode for Theorem 1.10 is presented in Algorithm 5. Following the structure of previous sections, we partition the proof of Theorem 1.10 into its privacy and utility proofs. We first present the privacy proof in Section 7.1.1 and then the approximation proof in Section 7.1.2. Finally, we bring the arguments together in Section 7.1.3 to complete the proof of Theorem 1.10.
7.1.1 Privacy Proof
Theorem 7.1.
Algorithm 5 is $\varepsilon$-node differentially private.
Proof.
Finally, for each of the $T$ rounds, we have that Algorithms 5, 5 and 5 can be analyzed as an instance of the Laplace mechanism with privacy parameter $\varepsilon^{\prime}$ and query $c_{0}$. In a given iteration, each bidder makes only a single bid so the sensitivity of the query is $1$, implying that this part is $\varepsilon^{\prime}$-differentially private. Applying the concurrent composition theorem, we have that Algorithms 5, 5 and 5 is $T\varepsilon^{\prime}$-differentially private. Finally, by concurrent composition, the entire algorithm is $T\varepsilon^{\prime}+T\varepsilon^{\prime}+T\varepsilon^{\prime}/\eta\leq\varepsilon$-differentially private. ∎
7.1.2 Utility Proof
@@
$(s-36c\log(n)/\varepsilon^{\prime})N-Q.$
$N\leq(\text{OPT}+Q)/(s-36c\log(n)/\varepsilon^{\prime}).$
Thus, the total number of bidders who were unmatched is at most
$B\geq T\cdot(Q-36c\log(n)/\varepsilon^{\prime}).$ (1)
Combining our two estimates in Eqs. 1 and 2, we have
implying that
7.1.3 Putting it Together
Proof of Theorem 1.10.
$w_{\mu(v),v}-p_{\mu(v)}\geq w_{\mu^{*}(v),v}-p_{\mu^{*}(v)}-\eta.$
Next, observe that if a good $u$ has $p_{u}>0$, this means that at least $s-18c\log(n)/\varepsilon^{\prime}$ bidders were (temporarily) matched with it, due to Algorithm 5. This directly implies that the number of bidders matched with $u$ at the termination is at least $s-36c\log(n)/\varepsilon^{\prime}$, because goods only unmatch with bidders until they are matched with $s-36c\log(n)/\varepsilon^{\prime}$ bidders (Line 5). Thus, there can be at most $\text{OPT}/(s-36c\log(n)/\varepsilon^{\prime})$ goods with $p_{u}>0$. For each of these goods, we have $p^{*}_{u}\cdot N_{u}^{*}-p_{u}\cdot N_{u}\leq s$, so we have
8 Arbitrary Edge-Order Continual Release Implicit Matchings
@@
We now give algorithms for matchings in the continual release model. We give algorithms that satisfy edge-DP and node-DP in two different types of input streams. Specifically, we consider two types of insertion-only streams where nodes and edges can be inserted but not deleted: 1) arbitrary-order edge arrival streams and 2) adjacency-list order streams.
In this section, we focus on arbitrary-order edge arrival streams, where edge insertions arrive in an arbitrary order, and show that our edge and node-DP implicit matching algorithms can be implemented in this model. Our results for the second stream type (adjacency-list order streams) is presented in Section 9. To avoid redundancy, we present the edge-DP algorithm (Theorem 1.11), which is essentially a lazy-update version of our static algorithm (Algorithm 2), and then briefly discuss the minor changes needed to obtain a node-DP algorithm (Theorem 1.12). Our results uses a variant of the standard SVT tool that allows us to answer “above” $c$ times, which we recall in Section A.3 for completeness.
See 1.11
Implementing the node-DP arboricity sparsifier of Section 6.2.1 in the continual release model leads to Theorem 1.12, our main continual release node-DP result restated below.
See 1.12
8.1 Proof of Theorem 1.11
@@
We begin with the pseudocode of the $\varepsilon$-edge DP algorithm in Algorithm 6. Then, we prove privacy and utility separately in Sections 8.1.1 and 8.1.2, respectively. The proof of privacy is non-trivial since we cannot simply apply composition. Instead, we directly argue by the definition of privacy.
@@ 7             $j_{t}\leftarrow j_{t}+1$
10             $\operatorname{solution}\leftarrow$ ($\nicefrac{{\varepsilon}}{{3c}}$)-edge DP implicit $b$-matching using Algorithm 2 on input $G_{t}$ (Theorem 1.3)
11             $\operatorname{estimate}\leftarrow$ ($\nicefrac{{\varepsilon}}{{3c}}$)-edge DP estimate of the size of the current maximum matching
@@ 8.1.1 Privacy Proof
Now, conditioned on the event that $j_{t-1}=J_{t-1}$, it holds that $j_{t}$ is independent of $\operatorname{solution}_{t-1}$ and $\operatorname{estimate}_{t-1}$. Hence the first product is at most $e^{\varepsilon/3}$ by the privacy guarantees of the SVT (Theorem A.3). For the second and third products, remark that given $j_{t}=J_{t-1}=\tilde{j}_{t}$, then $\operatorname{solution}_{t}=B_{t-1}=\widetilde{\operatorname{solution}}_{t}$ with probability $1$ and similarly for $\operatorname{estimate}_{t}$. Since $(J_{t})_{t}$ is a non-decreasing non-negative integer sequence bounded above by $c=a\log(n)/\eta$, at most $c=a\log(n)/\eta$ of the ratios from the second and third products are not 1. We can bound each of these ratios by $e^{\varepsilon/3c}$ using the individual privacy guarantees of Algorithm 2 (Theorem 1.3) and the Laplace mechanism. This concludes the proof. ∎
8.1.2 Utility Proof